Operations Report – May 26th

SOL: 30May2018

Name of person filing report: Shannon Rupert

Non-nominal systems: Power System Control

Notes on non-nominal systems: The power system continues to
malfunction. Yesterday it totally crashed and I had to run the generator for several hours and then it stabilized. I was offsite today but understand the problems continued. It is critical as the URC starts tomorrow and with the television show filming next week that I find out what is wrong.

Generator (hours run): not exactly sure as I left at 5 am this morning. Yesterday I ran it for three hours in the afternoon and then turned it back on at dark.

Solar—64 % SOC

Diesel – not checked %

Propane – not checked psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –Not checked

Water (trailer) – 0 gallons.

Water (static) – not checked gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: not recorded

Toilet tank emptied: no

ATV’s Used: no

Oil Added? no

ATV Fuel Used: n/a

# Hours the ATVs were Used today: 0

Notes on ATVs: We have a new battery for the 300 but I have not had time to install it.

Deimos rover used: Yes


Beginning charge:

Ending charge:

Currently charging: no

Sojourner rover used: Yes


Beginning charge:

Ending charge:

Currently charging: no

Spirit rover used: no


Beginning charge:

Ending charge:

Currently charging: no

Opportunity rover used: no


Beginning charge:

Ending charge:

Currently charging: yes

Curiosity rover used: No


Beginning charge:

Ending charge:

Currently charging: No

HabCar used and why, where? Yes, for URC travel

General notes and comments: The power system issue is really upsetting and I don’t know what I am going to do about it.

Yesterday all the drain pipes underneath the Hab were replaced. There were a total od 4 greywater leaks and no blackwater leaks. The largest leak was the one on the shower drain and it was caused by melting of the pipe by the heat tape. Apparently the wrong heat tape had been used on the PVC pipe and we are very lucky we did not have a fire as a result of someone’s carelessness. Because of some issues with contractor personnel, we are not going to reskirt the Hab at this time. There simply isn’t enough people power to do it. Instead, since we removed the wood frame, I will finish raking everything out from under it, and it looks nice, like a lander should. This will also allow us to air it out over the summer, as I had originally planned. We also had a great bonfire last night with the wood.

Also because of contractor personnel issues, the swamp cooler was not installed today. I will find out tomorrow what is happening with it. We were not able to get a contractor to install am AC system for another three months, so the swamp cooler is a work around for the AC requirement for the TV show.

Only two pieces of the tunnel remain to be covered and they will have to wait until the URC is over.

Someone turned the heater on in the Hab today, apparently thinking it was an AC unit. I put a note over the thermostat so it won’t happen again.

Shelves were installed in the storage unit. This should help in the future.

The road is in very bad shape because it has been so dry and all the travel has created the sand pits much earlier this year. I will call the county tomorrow and see if they can work on the worse of the spots.

Summary of internet: Nothing to report

Summary of suits and radios: Suits will be moved to the RAM tomorrow morning. I couldn’t do it today as the RAM was filled with URC materials.

Summary of Hab operations: See above.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing at this time

Copyright © The Mars Society. All rights reserved. | Main Site